# Solving the Poisson equation using coupled Markov chains

### v3

@@ 4.4 Implementation and improvements
Tuning of unbiased MCMC. The proposed estimator relies on unbiased signed measures $\hat{\pi}$ of $\pi$, as in (3.5). In our experiments, we start by generating some lagged chains, with a lag $L=1$ by lack of a better guess, and record the meeting times $\tau$. We then re-define $L$ as a large quantile of the meeting times, and we set $k=L$. We choose $\ell$ as a multiple of $k$ such as $5k$, following suggestions in Jacob et al. (2020); Atchadé & Jacob (2024), to ensure a low proportion $k/\ell$ of discarded iterations. We then run unbiased MCMC with these choices of $L,k,\ell$. We expect the inefficiency of such unbiased MCMC estimators to be at least $\ell/(\ell-k)$ that of ergodic average MCMC, due to the proportion $k/\ell$ of discarded iterations.
Choice of $y$. The proposed estimator requires setting $y$ to define $g_{y}$ as in Definition 4 and its estimator in Definition 5. Then $G_{y}(x)$ is generated for various $x$ which are approximately distributed according to $\pi$. As $G_{y}(x)$ should preferably have a smaller cost and a smaller variance, we should set $y$ such that two chains starting at $x\sim\pi$ and $y$ are likely to meet quickly. Thus, $y$ should preferably be central with respect to $\pi$. We experiment with the choice of $y$ in Section 5.4.
Selection probabilities. To implement UPAVE we choose selection probabilities $\xi=(\xi_{1},\ldots,\xi_{N^{(j)}})$ given $N^{(j)}$, the number of atoms in the signed measure $\hat{\pi}^{(j)}$. We set these probabilities to $1/N^{(j)}$ as a default choice. We can also try to minimize the variance of the resulting estimators with respect to $\xi$. This requires information on the variance of $G$. Indeed, if we condition on the realizations of $\hat{\pi}^{(j)}$, $j\in\{1,2\}$, then the variance of the term
$\frac{\omega_{I^{(r,j)}}^{(j)}}{\xi_{I^{(r,j)}}^{(j)}}(h(Z_{I^{(r,j)}}^{(j)})-\hat{\pi}^{(i)}(h))G(Z_{I^{(r,j)}}^{(j)}),\;\text{where}\;I^{(r,j)}\sim\text{Categorical}(\xi_{1}^{(j)},\ldots,\xi_{N^{(j)}}^{(j)}),$ (4.5)
@@
$\alpha_{n}$ =$=\{\omega_{n}^{(j)}(h(Z_{n}^{(j)})-\hat{\pi}^{(i)}(h))\}^{2}\mathbb{E}\left[G(Z_{n}^{(j)})^{2}|Z_{n}^{(j)}\right],$
$\xi_{n}^{\star}$ $=\frac{\sqrt{\alpha_{n}}}{\sum_{n^{\prime}=1}^{N^{(j)}}\sqrt{\alpha_{n^{\prime}}}},\quad n=1,\ldots,N^{(j)}.$ (4.6)
The use of $\xi^{\star}$ leads to a second moment of (4.5) equal to $(\sum_{n=1}^{N^{(j)}}\sqrt{\alpha_{n}})^{2}$, but for any $\xi$ such that $\sum_{n=1}^{N^{(j)}}\xi_{n}=1$, the Cauchy–Schwarz inequality implies $\sum_{n=1}^{N^{(j)}}\alpha_{n}/\xi_{n}\geq(\sum_{n=1}^{N^{(j)}}\sqrt{\alpha_{n}})^{2}$. Therefore $\xi^{\star}$ in (4.6) results in the smallest variance of the term in (4.5). We experiment with the estimation of $\xi^{\star}$ in Appendix H.
Choice of $R$. We need to choose $R$, the integer such that $2R$ is the number of states at which the fishy function $g$ is estimated. We can guide the choice of $R$ numerically by monitoring the inefficiency defined as the product of expected cost and variance, which can be approximated from independent copies of UPAVE. In our experiments we observe gains in efficiency when setting $R$ to a value such that the cost of generating the two unbiased signed measures matches approximately the cost of $2R$ fishy function estimators. This way, at most half of the computing budget is allocated sub-optimally. We note that when we run UPAVE for a given choice of $R$, we can also easily output estimators corresponding to smaller values of $R$, at no extra cost, which helps in monitoring the effect of $R$.
Reservoir sampling. A naive implementation of UPAVE with Algorithm 4.1 could incur a large memory cost when each state in $\mathbb{X}$ is large, as in high-dimensional regression (Biswas et al., 2022), or phylogenetic inference (Kelly et al., 2023). Indeed, storing all the atoms of the generated signed measures might be cumbersome. However, for UPAVE we only need to select within each measure $R$ atoms at which to evaluate $h$ and to estimate $g$; see Line 3c in Algorithm 4.1. We can address the memory issue by setting $\xi_{k}^{(j)}=1/N^{(j)}$ for all $k$ and by using reservoir sampling (Vitter, 1985). This technique allows to sample $I^{(r,j)}$ uniformly in $\{1,\ldots,N^{(j)}\}$, $R$ times independently, without knowing $N^{(j)}$ in advance and keeping only $R$ objects in memory.
We mention other methodological variations that we do not investigate further in this manuscript. Instead of sampling $R$ atoms from each signed measure with replacement, we could sample without replacement. Also, the $2R$ estimators $G$ employed in (4.4) could be generated jointly instead of independently. In particular, we can couple $2R$ chains starting from $(Z_{I^{(r,j)}})_{r=1}^{R}$, $j\in\{1,2\}$ with a common chain starting from $y$: in other words, we could simulate a single coupling of $12R+$ 1 chains instead of simulating $2R$ couplings of two chains.
@@ 4.5 Multivariate extension
Using the Cramér–Wold theorem (Billingsley, 1995, Theorem 29.4), we can handle the case where the test function $h$ takes values in $\mathbb{R}^{d}$. Write $h(x)=(h_{1}(x),\ldots,h_{d}(x))$, and write $h_{0}=h-\pi(h)$. Consider the Poisson equation for each $h_{i}$ and introduce the associated solutions denoted by $g_{i}$, and $g=(g_{1},\ldots,g_{d})$. The sum $\sum_{s=0}^{t-1}(h_{i}(X_{s})-\pi(h_{i}))$ can be re-written as $\sum_{s=1}^{t-1}(g_{i}(X_{s})-Pg_{i}(X_{s-1}))+g_{i}(X_{0})-Pg_{i}(X_{t-1})$. Observe that $S_{i}=(g_{i}(X_{t})-Pg_{i}(X_{t-1}))_{t\geq 1}$ is a martingale difference sequence for which a central limit theorem applies, with asymptotic variance as in (1.3), with $g_{i}$ instead of $g$. Write $S=(S_{1},\ldots,S_{d})$. For any vector $t\in\mathbb{R}^{d}$, we find that $t^{T}S$ is a martingale difference sequence as well, and by the Cramér–Wold theorem the multivariate asymptotic variance is
@@
Next, the multivariate extension of the alternate representation in (4.1) is obtained by developing the product of terms, then by using $Pg=g-h_{0}$ pointwise, and elementwise. Eventually we obtain
@@ 5 Numerical experiments with asymptotic variance estimators
We implement the proposed estimators of $v(P,h)$ and investigate their distinctive features, first in an AR(1) example where $v(P,h)$ is analytically available. Then we consider three MCMC algorithms that are representative of methods commonly used in Bayesian data analysis, namely a Hamiltonian Monte Carlo algorithm targeting a posterior distribution in logistic regression (Heng & Jacob, 2019), a Gibbs sampler for Bayesian regression with shrinkage prior where the target is high-dimensional, multimodal and heavy-tailed (Biswas et al., 2022), and a particle marginal Metropolis–Hastings algorithm to infer parameters in state space models (Middleton et al., 2020). Appendix H provides further experiments in the Cauchy location model of Section 2.4, including a study of the effect of the selection probabilities $\xi$ in UPAVE.
In all examples below, we tune unbiased MCMC as suggested in Section 4.4, with $L$ obtained from preliminary runs, $k=L$ and $\ell=5k$, so that unbiased MCMC is at least 20% less efficient than ergodic average MCMC. We compare the proposed UPAVE with batch means (BM) and spectral variance (SV) estimators. We implement them as in the mcmcse package (Flegal et al., 2020), with details provided in Appendix D. These estimators leverage some of the latest developments in asymptotic variance estimation: 1) the batch size is selected automatically with the method of Liu et al. (2022), 2) the bias is reduced using the lugsail variants proposed in Vats & Flegal (2022), with lugsail parameter $r\in\{1,2,3\}$, and 3) each estimator is based on a number of parallel chains, as proposed in Agarwal & Vats (2022); Gupta & Vats (2020). We have also experimented with the spectrum0 function of the coda package (Plummer et al., 2006), which gave estimates with low bias but prohibitively large variance in the examples considered below, and the results are not reported.
@@ 5.1 AR(1)
We consider the autoregressive process $X_{t}=\phi X_{t-1}+W_{t}$, where $W_{t}\sim\text{Normal}(0,1)$, and $(W_{t})$ are independent. We set $\phi=0.99$. The initial distribution is $\pi_{0}=\text{Normal}(0,4^{2})$. The target distribution is $\text{Normal}(0,(1-\phi^{2})^{-1})$, and for $h:x\mapsto x$ the asymptotic variance is analytically available: $v(P,h)=(1-\phi)^{-2}=10^{4}$. We use a reflection-maximal coupling as described in Appendix A. Assumption 2 holds for all $\kappa>1$ by Proposition 11, taking $\lambda=1+\frac{\phi}{2}$, $V(x)=\left|x\right|+1$, $b=1+\sqrt{2/\pi}$ and $\bar{v}=\frac{2b}{1-\phi}$ therein. We also investigate quantitative bounds in Appendix E. We choose $k=500$, $L=500$, $\ell=5k$ for unbiased MCMC. The state $y\in\mathbb{X}=\mathbb{R}$, used to define $g=g_{y}$ in (2.3), is set as $y=0$. We can calculate that $g_{y}$ here is the function $x\mapsto(1-\phi)^{-1}x$.
@@
The performance of the proposed estimator of $v(P,h)$ is shown in Table 5.1. The columns correspond to: 1) $R$: the number of atoms in each measure at which the fishy function is estimated, 2) estimate: overall estimate of $v(P,h)$, obtained by averaging $M$ independent runs, 3) cost: average cost of each run, in units of MCMC transitions, 4) fishy cost: average cost associated with fishy function estimates within each run, 5) empirical variance of the proposed estimators, and 6) inefficiency defined as the product of variance and average cost (smaller is better). The results are based on $M=10^{3}$ independent replicates, and each entry shows a $95\%$ confidence interval obtained with the nonparametric bootstrap. We see that increasing $R$ increases the average cost but improves efficiency. Overall we obtain accurate estimates of $v(P,h)$ with parallel runs that each costs of the order of $10^{4}$ iterations.
Next we compare the proposed estimator with the following “naive” strategy: generate a chain of length $T$ (post burn-in), compute the estimate $T^{-1}\sum_{t=0}^{T-1}h(X_{t})$, repeat $M$ times independently and compute the empirical variance of the $M$ estimates. We set $T=13,000$ to match the cost of our estimator with $R=50$. We find that the naive strategy has a bias (as $M\to\infty$) equal to $-77$, and an asymptotic variance of $2\times 10^{8}$. This is about 10 times larger than the variance of the proposed estimator reported in Table 5.1.
@@
We continue the comparison with batch means and spectral variance estimators. We compute these estimators of $v(P,h)$ using chains of lengths in $\{10^{4},10^{5},10^{6}\}$, and numbers of parallel chains in $\{1,2,4,8\}$. We base the results in this section on $400$ independent trajectories of length $10^{6}$, so for example we obtain $400$ independent estimators based on one chain, $200$ based on two chains, etc. For each configuration we approximate the mean squared error (MSE), and the total cost is equal to number of chains multiplied by the time horizon. Finally, we report the MSE that would be achieved by our proposed method, using $R=50$, if we generated sequentially a number of independent replicates of UPAVE corresponding to the given total cost. The comparison here does not account for any potential speed up on parallel architectures. The results are shown in Figure 5.1, where both axes are on logarithmic scale. Overall we note that the proposed method is worse than standard asymptotic variance estimators when the total computing budget is low. However, we observe that the MSE of the different estimators converge at different rates as a function of the total cost, as predicted by the theory reviewed in Appendix D. The proposed estimator has a smaller MSE when the computing budget is large enough.
@@
Finally we produce similar plots for the bias instead of the mean squared error, shown in Figure 5.2. The figure includes only batch means and spectral variance estimators since the proposed method is unbiased by design. We see that the $r$ lugsail parameter has a strong effect on the bias. In particular, the value $r=1$ that resulted in the smallest MSE in Figure 5.1 corresponds to a noticeable negative bias.
@@ 5.2 Hamiltonian Monte Carlo for Bayesian logistic regression
For the logistic regression model with a fixed prior standard deviation $\sigma_{a}>0$, the target is strongly log-concave, and the gradient of the log density is Lipschitz (see e.g. supplementary materials of Heng & Jacob (2019)). Under these conditions, Theorem 2 of Heng & Jacob (2019) states that their proposed coupling of a variant of Hamiltonian Monte Carlo is shown to lead to meeting times with Geometric tails, if the stepsize parameter and the number of leapfrog steps per iteration are small enough. However, the arguments in Heng & Jacob (2019) do not hold when $\sigma_{a}$ is allowed to be arbitrarily close to zero. In that case, the formal study of the meeting times is an open question. We employ a coupling of Hamiltonian Monte Carlo described in Appendix F. To sample the momentum variables we use a diagonal mass matrix, with entries approximately equal to the inverses of the posterior variances from 2000 samples generated by Stan (Carpenter et al., 2017). We use a number of leapfrog steps uniformly distributed between $1$ and $20$ at each iteration, and the stepsize is Exponential with rate $5$. The coupling strategy involves reflection-maximal couplings when the number of leapfrog steps is equal to one, and common random numbers otherwise. We compare HMC with the Metropolis-adjusted Langevin Algorithm (MALA), with the same random stepsize; we do not try to optimize the stepsize for either algorithm. The HMC algorithm employs here on average ten times more gradient evaluations than MALA per iteration. Yet in our implementation the elapsed real time of an HMC iteration is less than twice that of a MALA iteration.
Figure 5.3 shows three independent HMC trajectories (left), a histogram of the parameter of interest $\beta_{c}$ (middle), and upper bounds on the total variation (TV) distance to stationarity for HMC and for MALA (see Appendix C). From the plot we set $k=L=500,\ell=5k$ for unbiased HMC and $k=L=10^{4},\ell=5k$ for unbiased MALA. To implement UPAVE we need to define $g=g_{y}$. We draw $y$ (once and for all) from a Normal approximation of the posterior, obtained from 2000 samples from Stan.

### v4

@@ 4.4 Implementation and improvements
Tuning of unbiased MCMC. The proposed estimator relies on unbiased signed measures $\hat{\pi}$ of $\pi$, as in (3.5). In our experiments, we start by generating some lagged chains, with a lag $L=1$ by lack of a better guess, and record the meeting times $\tau$. We then re-define $L$ as a large quantile of the meeting times, and we set $k=L$. We choose $\ell$ as a multiple of $k$ such as $5k$, following suggestions in Jacob et al. (2020), to ensure a low proportion $k/\ell$ of discarded iterations. We expect the inefficiency of the resulting unbiased MCMC estimators to be at least $\ell/(\ell-k)$ that of ergodic average MCMC, due to the proportion $k/\ell$ of discarded iterations.
Choice of $y$. The proposed estimator requires setting $y$ to define $g_{y}$ as in Definition 4 and its estimator in Definition 5. Then $G_{y}(x)$ is generated for various $x$ which are approximately distributed according to $\pi$. As $G_{y}(x)$ should preferably have a smaller cost and a smaller variance, we should set $y$ such that two chains starting at $x\sim\pi$ and $y$ are likely to meet quickly. Thus, $y$ should preferably be central with respect to $\pi$.
Selection probabilities. To implement UPAVE we need to choose selection probabilities $\xi=(\xi_{1},\ldots,\xi_{N})$ given $N$, the number of atoms in the signed measure $\hat{\pi}$. We set these probabilities to $1/N$ as a default choice. We can also try to minimize the variance of the resulting estimators with respect to $\xi$. This requires information on the variance of $G$. Indeed, if we condition on the realization of $\hat{\pi}$, then the variance of the term
$\frac{\omega_{I_{r}}}{\xi_{I_{r}}}(h(Z_{I_{r}})-\hat{\pi}(h))G(Z_{I_{r}}),\;\text{where}\;I_{r}\sim\text{Categorical}(\xi_{1},\ldots,\xi_{N}),$ (4.5)
@@
$\alpha_{n}$ =$=\{\omega_{n}(h(Z_{n})-\hat{\pi}(h))\}^{2}\mathbb{E}\left[G(Z_{n})^{2}|Z_{n}\right],$
$\xi_{n}^{\star}$ $=\frac{\sqrt{\alpha_{n}}}{\sum_{n^{\prime}=1}^{N}\sqrt{\alpha_{n^{\prime}}}},\quad n=1,\ldots,N.$ (4.6)
The use of $\xi^{\star}$ leads to a second moment of (4.5) equal to $(\sum_{n=1}^{N}\sqrt{\alpha_{n}})^{2}$, but for any $\xi$ such that $\sum_{n=1}^{N}\xi_{n}=1$, the Cauchy–Schwarz inequality implies $\sum_{n=1}^{N}\alpha_{n}/\xi_{n}\geq(\sum_{n=1}^{N}\sqrt{\alpha_{n}})^{2}$. Therefore, $\xi^{\star}$ in (4.6) results in the smallest variance of the term in (4.5). We experiment with the estimation of $\xi^{\star}$ in Appendix H.
Choice of $R$. We need to choose $R$, the integer such that $2R$ is the number of states at which the fishy function $g$ is estimated. We can guide the choice of $R$ numerically by monitoring the inefficiency defined as the product of expected cost and variance, which can be approximated from independent copies of UPAVE. In our experiments we observe gains in efficiency when setting $R$ to a value such that the cost of generating the two unbiased signed measures matches approximately the cost of $2R$ fishy function estimators. This way, at most half of the computing budget is allocated suboptimally. We note that when we run UPAVE for a given choice of $R$, we can also easily output estimators corresponding to smaller values of $R$, at no extra cost, which helps in monitoring the effect of $R$.
Reservoir sampling. A naive implementation of UPAVE with Algorithm 4.1 could incur a large memory cost when each state in $\mathbb{X}$ is large, as in high-dimensional regression (Biswas et al., 2022), or phylogenetic inference (Kelly et al., 2023). Indeed, storing all the atoms of the generated signed measures might be cumbersome. However, for UPAVE we only need to select within each measure $R$ atoms at which to evaluate $h$ and to estimate $g$; see Line 3c in Algorithm 4.1. We can address the memory issue by setting $\xi_{n}=1/N$ for all $n$ and by using reservoir sampling (Vitter, 1985). This technique allows sampling $I_{r}$ uniformly in $\{1,\ldots,N\}$, $R$ times independently, without knowing $N$ in advance and keeping only $R$ objects in memory.
We mention other methodological variations that we do not investigate further in this manuscript. Instead of sampling $R$ atoms from each signed measure with replacement, we could sample without replacement. Also, the $2R$ estimators $G$ employed in (4.4) could be generated jointly instead of independently. In particular, we can couple $2R$ chains starting from $(Z_{I^{(j)}_{r}})_{r=1}^{R}$, $j\in\{1,2\}$ with a common chain starting from $y$: in other words, we could simulate a single coupling of $12R+$ 1 chains instead of simulating $2R$ couplings of two chains.
@@ 4.5 Multivariate extension
Using the Cramér–Wold theorem (Billingsley, 1995, Theorem 29.4), we can handle test functions $h$ taking values in $\mathbb{R}^{d}$: $h(x)=(h_{1}(x),\ldots,h_{d}(x))$. Write $h_{0}=h-\pi(h)$. Consider the Poisson equation for each $h_{i}$ and introduce the associated solutions denoted by $g_{i}$, and $g=(g_{1},\ldots,g_{d})$. The sum $\sum_{s=0}^{t-1}(h_{i}(X_{s})-\pi(h_{i}))$ can be re-written as $\sum_{s=1}^{t-1}(g_{i}(X_{s})-Pg_{i}(X_{s-1}))+g_{i}(X_{0})-Pg_{i}(X_{t-1})$. Observe that $S_{i}=(g_{i}(X_{t})-Pg_{i}(X_{t-1}))_{t\geq 1}$ is a martingale difference sequence for which a central limit theorem applies, with asymptotic variance as in (1.3), with $g_{i}$ instead of $g$. Write $S=(S_{1},\ldots,S_{d})$. For any vector $t\in\mathbb{R}^{d}$, we find that $t^{T}S$ is a martingale difference sequence as well, and by the Cramér–Wold theorem the multivariate asymptotic variance is
@@
Next, the multivariate extension of the alternate representation in (4.1) is obtained by developing the product, then by using $Pg=g-h_{0}$ pointwise and elementwise. We obtain
@@ 5 Numerical experiments with asymptotic variance estimators
We implement the proposed estimators of $v(P,h)$ and investigate their distinctive features, first in an AR(1) example where $v(P,h)$ is analytically available. Then we consider MCMC algorithms that are representative of methods commonly used in Bayesian data analysis. In Section 5.2, we consider a Hamiltonian Monte Carlo algorithm targeting a posterior distribution in logistic regression with random effects (Heng & Jacob, 2019). Supplementary material (Douc et al., 2024) includes experiments with a Gibbs sampler for Bayesian regression with shrinkage prior where the target is high-dimensional, multimodal and heavy-tailed (Biswas et al., 2022), and a particle marginal Metropolis–Hastings algorithm to infer parameters in state space models (Middleton et al., 2020).
In all examples below, we tune unbiased MCMC as suggested in Section 4.4, with $L$ obtained from preliminary runs, $k=L$ and $\ell=5k$, so that unbiased MCMC is at least 20% less efficient than ergodic average MCMC. We compare the proposed UPAVE with batch means (BM) and spectral variance (SV) estimators. We implement them as in the mcmcse package (Flegal et al., 2020), with details provided in Appendix C. In particular, the batch size is selected automatically with the method of Liu et al. (2022), the bias is reduced using the lugsail variants proposed in Vats & Flegal (2022), with lugsail parameter $r\in\{1,2,3\}$, and each estimator is based parallel chains, as proposed in Agarwal & Vats (2022); Gupta & Vats (2020). We have also experimented with the spectrum0 function of the coda package (Plummer et al., 2006), which gave estimates with low bias but prohibitively large variance in the examples considered below, and the results are not reported.
@@ 5.1 AR(1)
We consider the autoregressive process $X_{t}=\phi X_{t-1}+W_{t}$, where $W_{t}\sim\text{Normal}(0,1)$, and $(W_{t})$ are independent. We set $\phi=0.99$. The initial distribution is $\pi_{0}=\text{Normal}(0,4^{2})$. The target distribution is $\text{Normal}(0,(1-\phi^{2})^{-1})$, and for $h:x\mapsto x$ the asymptotic variance is analytically available: $v(P,h)=(1-\phi)^{-2}=10^{4}$. We use a reflection-maximal coupling as described in Appendix A.1. Assumption 2 holds for all $\kappa>1$ by Proposition 11, taking $\lambda=1+\frac{\phi}{2}$, $V(x)=\left|x\right|+1$, $b=1+\sqrt{2/\pi}$ and $\bar{v}=\frac{2b}{1-\phi}$ therein. We also investigate quantitative bounds in Appendix D. We choose $k=500$, $L=500$, $\ell=5k$ for unbiased MCMC. The state $y\in\mathbb{X}=\mathbb{R}$, used to define $g=g_{y}$ in (2.3), is set as $y=0$. We can calculate that $g_{y}$ here is the function $x\mapsto(1-\phi)^{-1}x$.
@@
The performance of the proposed estimator of $v(P,h)$ is shown in Table 5.1. The columns correspond to: 1) $R$: the number of atoms in each measure at which the fishy function is estimated, 2) estimate: overall estimate of $v(P,h)$, obtained by averaging $M$ independent runs, 3) cost: average cost of each run, in units of MCMC transitions, 4) fishy cost: average cost associated with fishy function estimates within each run, 5) empirical variance of the proposed estimators, and 6) inefficiency defined as the product of variance and average cost (smaller is better). The results are based on $M=10^{3}$ independent replicates, and each entry shows a $95\%$ confidence interval obtained with the nonparametric bootstrap. Increasing $R$ leads to a higher average cost but better efficiency. Overall we obtain accurate estimates of $v(P,h)$ with parallel runs that each costs of the order of $10^{4}$ iterations.
Next we compare the proposed estimator with the following naive strategy: generate a chain of length $T$ (post burn-in), compute the estimate $T^{-1}\sum_{t=0}^{T-1}h(X_{t})$, repeat $M$ times independently and compute the empirical variance of the $M$ estimates. We set $T=13,000$ to match the cost of our estimator with $R=50$. We find that the naive strategy has a bias (as $M\to\infty$) equal to $-77$, and an asymptotic variance of $2\times 10^{8}$. This is about 10 times larger than the variance of the proposed estimator reported in Table 5.1.
@@
We continue the comparison with batch means and spectral variance estimators. We compute these estimators of $v(P,h)$ using chains of lengths in $\{10^{4},10^{5},10^{6}\}$, and numbers of parallel chains in $\{1,2,4,8\}$. We base the results in this section on $400$ independent trajectories of length $10^{6}$, so for example we obtain $400$ independent estimators based on one chain, $200$ based on two chains, etc. For each configuration we approximate the mean squared error (MSE), and the total cost is equal to number of chains multiplied by the time horizon. Finally, we report the MSE that would be achieved by our proposed method, using $R=50$, if we generated sequentially a number of independent replicates of UPAVE corresponding to the given total cost. The comparison here does not account for any potential speed up on parallel architectures. The results are shown in Figure 5.1, where both axes are on logarithmic scale. The proposed method is worse than standard estimators when the total computing budget is low. However, we observe that the MSE of the different estimators converge at different rates, as predicted by the theory reviewed in Appendix C. The proposed estimator thus has a smaller MSE when the computing budget is large enough.
@@
Finally, we produce similar plots for the bias instead of the mean squared error, shown in Figure 5.2. The figure includes only batch means and spectral variance estimators since the proposed method is unbiased by design. We see that the $r$ lugsail parameter has a strong effect on the bias. In particular, the value $r=1$ that resulted in the smallest MSE in Figure 5.1 corresponds to a noticeable negative bias.
@@ 5.2 Hamiltonian Monte Carlo for Bayesian logistic regression
For the logistic regression model with a prior standard deviation $\sigma_{a}>0$, the target is strongly log-concave, and the gradient of the log density is Lipschitz (see e.g. supplementary materials of Heng & Jacob (2019)). Under these conditions, Theorem 2 of Heng & Jacob (2019) states that their coupling of a variant of Hamiltonian Monte Carlo leads to meeting times with Geometric tails, if the stepsize parameter and the number of leapfrog steps are small enough. However, the arguments in Heng & Jacob (2019) do not hold when $\sigma_{a}$ is allowed to be arbitrarily close to zero. In that case, the formal study of the meeting times is an open question. We employ a coupling of Hamiltonian Monte Carlo described in Appendix E. To sample the momentum variables we use a diagonal mass matrix, with entries approximately equal to the inverses of the posterior variances from 2000 samples generated by Stan (Carpenter et al., 2017). We use a number of leapfrog steps uniformly distributed between $1$ and $20$ at each iteration, and the stepsize is Exponential with rate $5$. The coupling strategy involves reflection-maximal couplings when the number of leapfrog steps is equal to one, and common random numbers otherwise. We compare HMC with the Metropolis-adjusted Langevin Algorithm (MALA), with the same random stepsize; we do not try to optimize the stepsize for either algorithm. The HMC algorithm employs here on average ten times more gradient evaluations than MALA per iteration. Yet in our implementation the elapsed real time of an HMC iteration is less than twice that of a MALA iteration.
Figure 5.3 shows three independent HMC trajectories (left), a histogram of the parameter of interest $\beta_{c}$ (middle), and upper bounds on the total variation (TV) distance to stationarity for HMC and for MALA (see Appendix A.3). From the plot we set $k=L=500,\ell=5k$ for unbiased HMC and $k=L=10^{4},\ell=5k$ for unbiased MALA. To implement UPAVE we need to define $g=g_{y}$. We draw $y$ (once and for all) from a Normal approximation of the posterior, obtained from 2000 samples from Stan (Carpenter et al., 2017).